(1 point) Perform the Gram-Schmidt process on the following sequence of vectors. 4 -3 5 8 0 -5 X= y = ,Z = 8 -3 -2
Answer #1

Q.Ansoy Ginen vectors: y za 5 1-5 8 if X are Recall it rectors then Step va se Az which is basis) of some set y = - <2₂, 6,>Step 10) 4 8 8 and 08 ų = y- M a: (१,१११ n then tu, et ८१,१) %3 -3 47 = 101 । 8 8 १+914 -3/ [4 8 8 -3 24 4 R 8 8 too cuin 8 Sand % = z - czu <z, Q> 4 - LE, QZų 2o, e> 4) CZ, 6, >=25 8 20-40-16= -36 4 <4,07 < 4 T6+64 +64 4 8 8 = 144 & Lz, 4 >=4 5 -5 -5 4 가 A Tdi 8 2 - 5+1-4 1-5+2 +4 L-2+2-2 2 Step 2 let 4 = I le he vou,= - 4 8 V4? +82 +82 4 8 8 12 8 . + u 3 2 and ų =ų Telland ų 2 ( 21) 24 4* | 2 (1) 3 - 1 hence then set 4,4 es a Orthonomer и . Ansis